OpenStudy (anonymous):

What is the specific heat of an unknown substance if 2000 J of energy are required to raise the temperature of 4 grams of the substance 5 degrees Celsius

OpenStudy (chmvijay):

Q=M S delta T S=Q/M*delta T Delta T= 5°C M= 4 gram Q= 2000J can u do now

OpenStudy (chmvijay):

just find S Specific heat of an unknown substance

OpenStudy (anonymous):

um okk thanks for ur help :/

OpenStudy (chmvijay):

ur welcome :)

OpenStudy (anonymous):

i have non idea what im doing :P

OpenStudy (chmvijay):

mean...

OpenStudy (anonymous):

ok i will try

OpenStudy (chmvijay):

S= 2000 / 5*4 =2000/20 =100 J/g °C =0.1 KJ/g °C